OpenStudy (anonymous):

write this quadratic equation in standard form. x squared +2x=15 please explain :)

OpenStudy (anonymous):

x^2-2x+1=0

OpenStudy (callisto):

both sides - 15 so that the 15 on the right hand side can be cancel and give you a 0 on the right

OpenStudy (anonymous):

could u like show da work please im helpless

OpenStudy (callisto):

x squared +2x=15 x squared +2x -15=15-15 x squared +2x -15=0

OpenStudy (anonymous):

 Now it's in standard form